• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 13:11:19 on July17,2003

    Entry number 104254

    This entry is a followup to: 103856

    keyword=L-arm S2 TDC spectrum

    Rob did some work to improve the L-arm S2 TDCs (problem first reported in entry #103856). We used run #21380 (the first L-arm run with carbon target installed) to check, the TDCs look much better now. S2 ADCs have much smaller signals but Rob said this is expected.

    FIGURE 1

    FIGURE 2